1. Iodic acid, HIO3, can be prepared by reacting iodine, I2, with concentrated nitric


acid. The skeleton equation is:


I2(s) + NO3


-


(aq) → IO3


-


(aq) + NO2(g). Balance this equation using the half reaction method.

1
Expert's answer
2022-06-06T21:48:31-0400

I2 + HNO3 = HIO3 + NO2 + H2O

On top of each element, we write the degree of oxidation:

I20 + [N5+O32-]- = [I5+O32-]- + N4+O22-


We write down the elements whose oxidation state has changed and determine how many electrons they have received or donated:

I2 - 10 e- => 2 I5+

N5+ + e- => N4+


We write the number of electrons given by I2 in front of N5 + and the number of electrons received by N5 + in front of I2 as a coefficient.


Answer:

I2 + 10 HNO3 = 2 HIO3 + 10 NO2 + 4H2O